# Soft Reasoning on Uncertain Knowledge Graphs

### v1

@@ 4.3 Two calibration strategies
Debiasing. The confidence function $\hat{P}$ of UKGE is biased towards zero. We propose a debiasing strategy for the inference. That is, we modify the soft requirements $\alpha$ as $\alpha-\Delta_{\alpha}$. We can see this simple debiasing strategy improves the performance. SRC with this strategy is denoted as SRC(D).
Learning. The pre-trained UKGE may not be optimal for SRC in the OWA. We propose the calibration by learning (L) strategy by learning the calibrated confidence function. Specifically, we calibrate the confidence function $\hat{P}$ by
@@
The parameterization of learnable affine transformation also follows the previous adaptation in Arakelyan et al. (2023).
@@
$\mathcal{L}=\sum_{U[\phi](s)>0}(\hat{U}[\phi,\hat{P}_{c}](s)-U[\phi](s))^{2},$ (54)
where $\hat{U}[\phi]$ represents the predicted utility vector of a soft query $\phi$. SRC with this strategy is denoted as SRC(L).
Importantly, we only need to train the calibration transformation in the cases of $\alpha=0$, achieving a simpler training strategy but better generalization capability when compared to the QE+NE baselines, as will be presented in Section 6.2.
5 Dataset Construction
Our proposed dataset targets the open world assumption for UKG and considers soft queries where the utility vector will be updated based on the underlying UKG. We utilize three standard UKGs with train/validation/test splits: CN15k (Chen et al., 2019) for encompassing commonsense, PPI5k (Chen et al., 2019) for biology, and O*NET20K (Pai & Costabello, 2021) for employment domains. We then provide a brief overview of the proposed dataset and the details of dataset construction can refer to Appendix G.
5.1 Queries and evaluation protocols
Compared with the observed uncertain KG, the underlying uncertain KG contains new facts that will update the utility vectors of some soft queries. Only these specific queries are considered to be meaningful and are included for evaluation. We adopt metrics from learning to rank as our evaluation protocol, including MAP, NDCG, Spearman’s rank correlation coefficient $\rho$, and Kendall’s rank correlation coefficient $\tau$.
We select 1P, 2P, 2I, 2IN, and 2IL as train query types and add 3IN, INP, IP, 2M, IM, 2U, and UP as validation/test query types. The training query types have included the basic logic operations and we aim to assess the ability of ML methods to generalize to unseen query types. We present the query structures of these query types in Figure 2.
@@ 5.2 Soft requirements
For the $\alpha$ parameter, we establish connections with the percentile value of the relation to represent confidence requirements effectively. We assign specific percentiles to different necessity levels: the 25th, 50th, and 75th percentiles correspond to “low”, “normal”, and “high” confidence requirements, respectively. We ensure that a “zero” requirement is assigned when the necessity requirement reaches 0.
For the $\beta$ importance setting, we employ two strategies: “equal” and “random”. Under the “equal” strategy, all requirements are assigned an importance value of $1.0$. In contrast, the “random” strategy introduces variability by assigning random decimal numbers between 0 and 1 to represent the importance of each soft constraint.
@@ 6 Experiments
In this section, we empirically explore how to answer soft queries. First, we compare our proposed method with two classical query embedding approaches. To enable query embedding to be aware of soft requirements, we modify the relation projection network with number embedding (NE), where the detailed information is provided in Appendix F. Additionally, we examine the impact of two parameters $\alpha$ and $\beta$ on both kinds of approaches. Finally, we evaluate the performance of large language models on soft queries with clear natural language semantics. The implementation details of these experiments are in Appendix C.
Table 2: Result of answering soft queries. Logic+NE and ConE+NE refer to the query embedding with number embedding extensions. SRC is our inference method, and SRC(D), SRC(L), and SRC(D+L) is explained in Section 4.3.
UKG Models $\tau$ AVG. $\rho$ AVG. MAP AVG. NDCG
1P 2P 2I 2IN 2IL 2M 2U 3IN IP IM INP UP AVG
@@ 6.1 Main results
Models analysis. The main result is shown in Table 2, indicating that our proposed method exhibits significantly higher performance compared to query embedding methods such as LogicE (Luus et al., 2021) and ConE (Zhang et al., 2021). Furthermore, by leveraging the two calibration strategies: debiasing(D) and learning(L), explained in Section 4.3, the performance of our method is further enhanced, achieving the best results across most graphs and metrics on average. The results demonstrate the superior performance of training on the case $\alpha=0$ explained in Section 4.3.
@@
Query types analysis. Although ConE performs well on some trained query types, it struggles when faced with newly emerged and more complex query types like INP and IM. In contrast, our method exhibits excellent performance across the majority of query types, demonstrating its robust combinatorial generalization capability. Our method particularly excels in handling challenging query types that involve existential variables, such as 2P, 2M, IM, and INP.
6.2 The impact of soft requirements
@@
The two parameters, $\alpha$ and $\beta$ play a crucial role in controlling soft constraints, thus we construct settings with varying values. Specifically, we select “zero”(Z) and “random”(R) for $\alpha$, and “equal”(E) and “random”(R) for $\beta$, as explained in Section 5.2. Moreover, we sample 12 query types from O*NET20k KG. The detailed construction is in Appendix H. For ConE, we train it on each setting and test it across all settings. As for SRC, we directly test it on all settings.
@@
Model Train Test AVG
@@
6.3 The comparison with large language model
We devise an evaluation framework to assess the performance of Large Language Models (LLMs), benchmarking their powerful knowledge reasoning abilities. To ensure fairness of the comparison, we consider the queries sampled from CN15k because of the commonsense semantics and we select four candidate answers for each query. These queries have also been manually filtered and labeled to ensure clearness and correctness. We describe the syntax and semantics of soft queries using natural language, prompting LLMs to select the most suitable answer. For details on this setting construction, please refer to Appendix I.
Table 4: The accuracy of manually annotated queries.
Model GPT-3.5-turbo GPT-4-preview SRC
Accuracy 34.3 37.8 48.9
The result is shown in Table 4, indicating that SRC achieves significantly higher accuracy compared to GPT-3.5-turbo and GPT-4-preview. Despite these queries being translated into natural language, LLMs still struggle to select the highest-scoring answer. This further highlights the difficulty of the soft queries we propose. Thus, continued development of neural-symbolic approaches is crucial, particularly in specialized domains such as employment (Pai & Costabello, 2021) and biology (Szklarczyk et al., 2023).
7 Conclusion
In this paper, we introduce a novel setting, soft queries on uncertain knowledge graph. The soft queries focus on the OWA for UKGs and require the incorporation of ML methods to estimate scores for new edges while handling semiring computations. Our proposed soft queries also consider the soft requirements to control the uncertainty of knowledge. To facilitate the research of soft queries, we construct a dataset consisting of three uncertain knowledge graphs. Furthermore, we propose a new neural-symbolic approach with both forward inference and backward calibration. Both theoretical analysis and experimental results demonstrate that our method has satisfactory performance.
8 Potential Impact
This paper focuses on soft queries on uncertain knowledge graphs and how to answer soft queries efficiently. The datasets we propose are derived from publicly available standard uncertain knowledge graphs and do not raise ethical concerns. However, it is important to acknowledge that our proposed method, SRC, has the potential to be utilized for commercial purposes, specifically in efficiently inferring private information from industrial-scale uncertain knowledge graphs. This aspect is an inherent consideration when proposing models to answer complex logical queries.
Appendix A Related Work
A.1 Complex logical queries
Answering complex logical queries over knowledge graphs is naturally extended from link prediction and aims to handle queries with complex conditions beyond simple link queries. This task gradually grows by extending the scope of complex logical queries, ranging from conjunctive queries (Hamilton et al., 2018) to Existential Positive First-Order (EPFO) queries (Ren et al., 2020), Existential First-Order (EFO) queries (Ren & Leskovec, 2020; Wang et al., 2021), real Existential First-Order queries (Yin et al., 2024). The primary method is query embedding, which maps queries and entities to a low-dimensional space. The form of embedding has been well investigated, such as vectors (Hamilton et al., 2018; Chen et al., 2022; Bai et al., 2022), geometric regions (Ren et al., 2020; Zhang et al., 2021), and probabilistic distributions (Ren & Leskovec, 2020; Choudhary et al., 2021; Yang et al., 2022; Wang et al., 2023a). These methods not only explore knowledge graphs embedding but also leverage neural logical operators to generate the embedding of complex logical queries.
There are also neural-symbolic models to answer complex logical queries. Amayuelas et al. (2021) and Arakelyan et al. (2023) employ gradient optimization techniques to estimate the embedding existential variables. Zhu et al. (2022) adapts graph neural network to execute relational projects and use logical operations over fuzzy sets to deal with more complex queries. Yin et al. (2024) and Bai et al. (2023) present efficient search algorithms based on link predictor over knowledge graphs. While symbolic methods demonstrate good performance and offer interpretability for intermediate variables, they often struggle to scale with larger graphs due to their high computational complexity.
Yin et al. (2023) present the challenge setting and introduce logical queries that involve multiple free variables. Many other models and datasets are proposed to enable answering queries with good performance and additional features, see Ren et al. (2023) for a comprehensive survey. However, to the best of our knowledge, there is currently no existing query framework specifically designed for uncertain knowledge graphs.
A.2 Uncertain knowledge graph embedding
The first line of research focuses on predicting the confidence score of uncertain relation facts. UKGE (Chen et al., 2019) was the pioneering effort to model triple plausibility as the activated product of these embedding vectors. UKGE incorporates soft probabilistic logic rules to provide the plausibility of unseen facts. Building upon this, BEUrRE (Chen et al., 2021a) utilizes complex geometric boxes with probabilistic semantics to represent entities and achieve better performance. Chen et al. (2021b) applies semi-supervised learning to predict the associated confidence scores of positive and negative samples. And Brockmann et al. (2022) and Liang (2023) both explore graph neural networks to represent and predict uncertain knowledge graphs.
The other line of research aims to address link prediction on uncertain knowledge graphs by fitting the likelihood of uncertain facts. To adjust the similar task, Pai & Costabello (2021) introduces FocusE, an additional layer to the knowledge graph embedding architecture. They provide variants of classical embedding methods such as TransE (Bordes et al., 2013), DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016).
Appendix B Constraint Satisfaction Problem and Soft Constraint Satisfaction Problems

### v2

@@ 4.3 Two calibration strategies
Debiasing. The confidence function $\hat{P}$ of UKGE is biased towards zero. We propose a debiasing strategy for the inference. We modify the soft requirements $\alpha$ as $\alpha-\Delta_{\alpha}$. We can see that this simple debiasing strategy improves performance. SRC with this strategy is denoted as SRC(D).
Learning. The pre-trained UKGE is not optimal for SRC in the incomplete uncertain KGs. We propose the calibration by learning (L) strategy by learning the calibrated confidence function. Specifically, we calibrate the confidence function $\hat{P}_{c}$ by learnable affine transformation (Arakelyan et al., 2023) as following:
@@
$[\rho_{\theta}(s,r,o),\lambda_{\theta}(s,r,o)]=\sum_{j\in\{s,r,o\}}(W_{j}\mathbf{e}_{j}+b_{j}).$
Here, $\rho_{\theta}(s,r,o)$ and $\lambda_{\theta}(s,r,o)$ are the affine parameters. $\mathbf{e}_{j}\in\mathbb{R}^{d}$ represents the embedding for entity or relation $j\in\{s,r,o\}$ with embedding dimension $d$. Meanwhile, $W_{j}\in\mathbb{R}^{2\times d}$ and $b_{j}\in\mathbb{R}^{2}$ are the learnable parameters associated with $j$.
@@
$\mathcal{L}=\sum_{\mathbf{u}(s)>0}(U(\phi(s),\hat{P}_{c})-U(\phi(s),\mathcal{P}))^{2},$ (44)
where $U(\phi(s),\hat{P}_{c})$ represents the predicted utility vector of a soft query $\phi$ according to Definition 6. SRC with this strategy is denoted as SRC(L).
Notably, we only need to train the calibration transformation in the cases of $\alpha=0$, achieving a simpler training strategy but better generalization capability when compared to the QE+NE baselines, as will be presented in Appendix B.
5 SQUK Dataset Construction
We provide a brief overview of dataset construction and the details can refer to Appendix K.
5.1 Useful queries and evaluation protocols
The validation/test uncertain knowledge graph incorporates new facts that will update the utility vectors of specific soft queries. Only these particular queries are considered meaningful and included in the evaluation. The evaluation of soft queries not only considers recall but also accounts for the values of the recalled answers. Therefore, we adopt metrics from the learning-to-rank framework Liu et al. (2009) as our evaluation protocol, which includes Mean Average Precision (MAP), Normalized Discounted Cumulative Gain (NDCG), Spearman’s rank correlation coefficient ($\rho$), and Kendall’s rank correlation coefficient ($\tau$).
We choose 1P, 2P, 2I, 2IN, and 2IL as train query types and add 3IN, INP, IP, 2M, and IM as validation/test query types. The training query types encompass basic operations, allowing us to evaluate the ability of machine learning methods to generalize to commonly used unseen query types (Yin et al., 2024). We visualize the structure of these types in Figure 4. Additionally, Table 8 shows the statistics of training queries of our dataset.
Table 2: The statistics of train queries.
KG 1P 2P 2I 2IN 2IL
PPI5k 9,724 9,750 9,754 1,500 1,500
O*NET20k 18,266 18,300 18,300 1,850 1,850
CN15k 52,887 52,900 52,900 5,300 5,300
5.2 Uncertain KGs
We utilize three standard uncertain KGs: CN15k (Chen et al., 2019) for encompassing commonsense, PPI5k (Chen et al., 2019) for biology, and O*NET20K (Pai and Costabello, 2021) for employment domains. These uncertain KGs are noisy and incomplete, requiring the ML models to predict the confidence values.
@@ 5.3 Soft requirements
For the $\alpha$ parameter, we establish connections with the percentile value of the relation to represent the necessity value effectively. We assign specific percentiles to different necessity levels: the 25th, 50th, and 75th percentiles correspond to “low”, “normal”, and “high” necessity criteria, respectively. We ensure that a “zero” requirement is assigned when the necessity criteria reaches 0. We also introduce a hybrid strategy that randomly selects necessity values, enabling a comprehensive evaluation.
For the $\beta$ importance setting, we employ two strategies: “equal” and “random”. Under the “equal” strategy, all importance values are assigned an importance value of $1.0$. In contrast, the “random” strategy introduces variability by assigning random decimal numbers between 0 and 1 to represent the importance of each soft atomic formula.
@@ 6 Experiments
In this section, we empirically explore how to answer soft queries. We mainly compare our method with generalized SoTA CQA models on SQUK dataset, including commonly used query embedding models and advanced symbolic search methods. Additionally, we evaluate the performance of advanced commercial LLMs on soft queries with clear natural language descriptions. The implementation details of these experiments are in Appendix G. We also conduct the ablation study regarding the distribution and impact of two parameters $\alpha$ and $\beta$ on both kinds of approaches, which is presented in Appendix B due to page limitations. We also provide the qualitative analysis of the experiments in Appendix C.
Table 3: Result of answering soft queries. Logic+NE and ConE+NE refer to the query embedding with number embedding extensions. SRC is our inference method, and SRC(D), SRC(L), and SRC(D+L) are explained in Section 4.3. The four metrics are all higher, indicating better performance.
Uncertain KG Models $\tau$ AVG. $\rho$ AVG. MAP AVG. NDCG
1P 2P 2I 2IN 2IL 2M 2U 3IN IP IM INP UP AVG.
@@ 6.1 Main results
Baselines We select two mainstream CQA methods as baselines: query embedding and symbolic search. Specifically, we focus on two classical query embedding methods: LogicE (Luus et al., 2021) and ConE (Zhang et al., 2021). To enable soft queries, we incorporate the relation projection network with Number Embedding (NE) and adjust the loss function accordingly.  [^4] Detailed information can be found in Appendix J. The forward inference of our method, SRC, is directly generalized from SoTA symbolic methods FIT (Yin et al., 2024), which serve as baselines for search methods.
@@
Models analysis. The main results are presented in Table 3, demonstrating that our proposed method significantly outperforms both query embedding methods and directly generalized symbolic methods. By leveraging the two calibration strategies, debiasing (D) and learning (L), as explained in Section 4.3, our method achieves superior results across most KGs and metrics on average.
@@
Query structure analysis. Although ConE performs well on some trained query types, it struggles with newly emerged query types and those involving negation, such as INP and IM. In contrast, our method exhibits excellent performance across the majority of query types, demonstrating robust combinatorial generalization capabilities on complex queries. Our method particularly excels in handling challenging query types that involve existential variables, such as 2P, 2M, IM, and INP, highlighting its advantages in these scenarios.
6.2 The comparison with LLMs
Table 4: The accuracy of manually annotated queries.
Model Llama3.3 70B Gemini-1.5-pro GPT-3.5-turbo GPT-4-preview SRC
Accuracy 42.6 37.1 34.3 37.8 48.9
We devise an evaluation framework to assess the performance of LLMs, benchmarking their powerful reasoning abilities over uncertain knowledge. To ensure fairness of the comparison, we consider the queries sampled from CN15k and we choose four candidate answers for each query. These queries have also been manually filtered and labeled to ensure clearness and correctness. We describe the syntax and semantics of soft queries using natural language, prompting LLMs to select the most suitable answer. The details on this setting construction can refer to Appendix M.
The results, shown in Table 4, indicate that even the simple symbolic SRC achieves significantly higher accuracy compared to Llama3.3 70B, Gemini-1.5-pro, GPT-3.5-turbo, and GPT-4-preview. This demonstrates that large language models (LLMs) struggle with complex arithmetic operations involving uncertain values of knowledge. Our evaluation is fair, as the required uncertain knowledge is derived from well-known commonsense KGs ConceptNet, and the logical operations are expressed in natural language. Nevertheless, even advanced commercial LLMs struggle to select the highest-scoring answer. This further emphasizes the difficulties presented by the proposed soft queries and highlights the ongoing need for the development of symbolic approaches.
7 Conclusion
In this paper, we introduce a novel setting, soft queries on uncertain KGs, which further extends complex logical queries on KG. The soft queries consider the incompleteness of large-scale uncertain KGs and require the incorporation of ML methods to estimate scores for new relational linking while handling semiring algebraic structures. Our proposed soft queries also propose the soft requirements inspired by soft constraint satisfaction problems to control the uncertainty of knowledge. To facilitate the research of soft queries, we construct a soft query answering dataset consisting of three uncertain KGs. Furthermore, we propose a new neural-symbolic approach with both forward inference and backward calibration. Both theoretical analysis and experimental results demonstrate that our method has satisfactory performance.
8 Limitation
Soft queries extend complex logical queries over Knowledge Graphs (KGs) by incorporating soft requirements within uncertain KGs. However, the scope of the proposed soft queries is limited, as it primarily focuses on conjunctive queries. While conjunctive queries form the foundation of complex logical queries, this restriction may hinder the expressiveness and applicability of the proposed soft queries. Furthermore, the dataset does not include cyclic queries, which are NP-complete, even though their complexity can be more easily addressed.
9 Potential Impact
Soft queries have the potential to perpetuate existing biases present in the underlying knowledge graphs. If these graphs contain skewed or discriminatory information, the results generated by soft queries may reflect and amplify these biases, leading to unfair outcomes in applications such as hiring, credit scoring, or law enforcement. This raises significant ethical concerns about fairness, as marginalized groups may be disproportionately affected by biased query results, resulting in systemic inequality.
The utilization of soft queries to extract information from knowledge graphs can pose serious privacy risks. If queries access sensitive personal data without proper safeguards, there is a potential for unauthorized disclosures that violate individuals’ privacy rights. This concern is heightened in contexts where the data might be used for profiling or surveillance, making it imperative to establish robust privacy protections and ethical guidelines to ensure that individuals’ information is handled responsibly and transparently.
Appendix
Appendix A Related Work
A.1 Complex logical queries
Answering complex logical queries over knowledge graphs is naturally extended from link prediction and aims to handle queries with complex conditions beyond simple link queries. This task gradually grows by extending the scope of complex logical queries, ranging from conjunctive queries Hamilton et al. (2018) to Existential Positive First-Order (EPFO) queries Ren et al. (2020), Existential First-Order (EFO) queries Ren and Leskovec (2020), real Existential First-Order queries Yin et al. (2024). The primary method is query embedding, which maps queries and entities to a low-dimensional space. The form of embedding has been well investigated, such as vectors Hamilton et al. (2018); Chen et al. (2022); Bai et al. (2022), geometric regions Ren et al. (2020); Zhang et al. (2021), and probabilistic distributions Ren and Leskovec (2020); Choudhary et al. (2021); Yang et al. (2022a); Wang et al. (2023a). These methods not only explore knowledge graphs embedding but also leverage neural logical operators to generate the embedding of complex logical queries.
There are also neural-symbolic models to answer complex logical queries. Gradient optimization techniques were employed to estimate the embedding existential variables Amayuelas et al. (2021); Arakelyan et al. (2023). Graph neural network Zhu et al. (2022) was adapted to execute relational projects and use logical operations over fuzzy sets to deal with more complex queries. Efficient search algorithms based on link predictor over knowledge graphs were presented Yin et al. (2024); Bai et al. (2023). While symbolic methods demonstrate good performance and offer interpretability for intermediate variables, they often struggle to scale with larger graphs due to their high computational complexity.
Many other models and datasets are proposed to enable answering queries with good performance and additional features, see the comprehensive survey Ren et al. (2023). However, to the best of our knowledge, there is currently no existing query framework specifically designed for uncertain knowledge graphs.
A.2 Uncertain knowledge graph embedding
The first line of research focuses on predicting the confidence score of uncertain relation facts. UKGE (Chen et al., 2019) was the pioneering effort to model triple plausibility as the activated product of these embedding vectors. UKGE incorporates soft probabilistic logic rules to provide the plausibility of unseen facts. Building upon this, BEUrRE (Chen et al., 2021a) utilizes complex geometric boxes with probabilistic semantics to represent entities and achieve better performance. Semi-supervised learning was applied Chen et al. (2021b)to predict the associated confidence scores of positive and negative samples. And Graph neural networks were used DBLP:journals/sigkdd/BrockmannKB22; Liang (2023) to represent and predict uncertain knowledge graphs.
The other line of research aims to address link prediction on uncertain knowledge graphs by fitting the likelihood of uncertain facts. To adjust the similar task, FocusE Pai and Costabello (2021) was introduced, an additional layer to the knowledge graph embedding architecture. They provide variants of classical embedding methods such as TransE (Bordes et al., 2013), DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016).
Appendix B Ablation study: The impact of soft requirements
@@
The two parameters, $\alpha$ and $\beta$ play a crucial role in controlling soft constraints, thus we construct settings with varying values. Specifically, we select “zero”(Z) and “random”(R) for $\alpha$, and “equal”(E) and “random”(R) for $\beta$, as explained in Section 5.3. Moreover, we sample 12 query types from O*NET20k KG. The detailed construction is in Appendix L. For ConE, we train it on each setting and test it across all settings. As for SRC, we directly test it on all settings.
@@
Model Train Test AVG.
@@
Appendix C Qualitative analysis
We select the following query for the LLM to conduct a qualitative analysis: “ Soft query: (bread, is related to, $f_{1}$, 0.3, 0.9) $\land$ (rice, is related to, $f_{1}$, 0.7, 0.3). Four candidate entities: basic, wheat, white, and food. ” Among the four candidates, basic and wheat are weakly related to the two soft constraints. While white and food exhibit higher relevance, the constraints related to bread carry greater importance. Consequently, the final answer is bread. However, some LLMs may often select rice due to a tendency to hallucinate the concept of white bread. Such errors highlight the challenges in accurately interpreting soft constraints and the need for more robust mechanisms to handle nuanced queries.
Appendix D Logical queries on knowledge graphs
Definition 9 (Knowledge graphs).
Let $\mathcal{E}$ be the set of entities and $\mathcal{R}$ be the set of relations. A knowledge graph is a set of triples $\mathcal{G}=\{(s_{i},r_{i},o_{i})\}$, where $s_{i},o_{i}\in\mathcal{E}$ are entities and $r_{i}\in\mathcal{E}$ is relation.
The fundamental challenge of knowledge graphs lies in dealing with the Open World Assumption (OWA). Unlike the Closed World Assumption (CWA), which considers only observed triples as facts, OWA acknowledges that unobserved triples may also be valid.
The study of logical queries on KG considers the Existential First-Order (EFO) queries, usually with one free variable Ren and Leskovec (2020); Wang et al. (2021); Yin et al. (2024).
Definition 10 (Syntax of existential first-order queries).
The disjunctive normal form of an existential first-order query $\Gamma$ is:
$\Gamma(y)=\gamma_{1}(y)\lor\cdots\lor\gamma_{q}(y),$ (45)
Definition 11.
$\Gamma(s/y)$ denotes the substitution of the entity $s$ for the variable $y$.
When all free variables are substituted, a query $\Gamma(y)$ is transformed into a sentence $\Gamma(s/y)$. Given $\mathcal{G}$, answering a query $\Gamma(y)$ means finding all substitutions, such that the sentence $\Gamma(s/y)$ is entailed by $\mathcal{G}$, i.e., $\mathcal{G}\models\Gamma(s/y)$. The answer set is defined as
Definition 12 (The Answer Set of first-order 1uery).
The answer set of an first-order query is defined by
$\mathcal{A}[\Phi(y)]=\{a\in\mathcal{E}|\text{ }\Phi(a)\text{ is True}\}$ (47)
The answers in the answer set derived from $\mathcal{G}_{\rm train}$ is easy answers. hard answers are the answers in the set difference between the answers from $\mathcal{G}_{\text{valid}}$ and $\mathcal{G}_{\text{train}}$ Wang et al. (2021); Ren et al. (2023). The traditional graph-matching methods can not find the answers introduced by new facts (Riesen et al., 2010). Thus, we should develop new methods.
D.1 ML-based method for logical queries on KG
Recent works are dedicated to introducing ML methods, i.e., knowledge graph embeddings, to generalize from $\mathcal{G}$ to $\widehat{\mathcal{G}}$, so that it approximates $\mathcal{G}_{\rm test}$.
Query Embeddings (QE). Query embedding methods generally map the query $\Gamma$ as an embedding (Ren and Leskovec, 2020; Liu et al., 2021; Wang et al., 2023b). One dominant way is to “translate” the procedure of solving logical formulas in Equation (45) into the set operations, such as set projection, intersection, union, and complement. Then, the neural networks are designed to model such set operations in the embedding space. We can see that direct modeling of set operations is incompatible with the concept of necessity and importance introduced in Section 3.
Inference methods. Other methods solve the open world query answering methods in a two-step approach (Arakelyan et al., 2021; Bai et al., 2023; Yin et al., 2024). In the first step, the pre-trained knowledge graph embedding estimates the $\mathcal{G}_{\rm test}$. Then, the answers are derived by the fuzzy logic inference or optimization. These methods rely on the standard logic calculation and cannot be directly applied to our SQUK setting introduced in Section 3.
Appendix E Connection with Open World Assumption
Evaluating queries over deductive question-answering systems generally follows either the closed-world assumption (CWA) or the open-world assumption (OWA) Reiter (1981). Under CWA, only known facts are considered true, whereas OWA assumes that the absence of knowledge does not imply falsity Reiter (1986). Since knowledge graphs (KGs) are often incomplete, knowledge graph completion (KGC) has been proposed to address this challenge. In KGC, the observed knowledge in KGs is divided into three nested subsets: $\mathcal{G}_{\text{train}}\subset\mathcal{G}_{\text{valid}}\subset\mathcal{G}_{\text{test}}$. The KGC model is trained on the $G_{\text{train}}$ subset and evaluated on the $G_{\text{valid}}$ subset to assess its performance. For complex logical query answering, the evaluation considers "hard answers," defined as the set difference between the answers from $G_{\text{valid}}$ and $G_{\text{train}}$ Wang et al. (2021); Ren et al. (2023). This approach assesses the model’s ability to handle incomplete knowledge and make inferences beyond the observed facts in the training set. The evaluation of complex logical query answering extends beyond CWA but does not fully align with OWA. The same applies to the evaluation of soft query answering. Evaluating under the Open-World Assumption represents a promising avenue for future work Yang et al. (2022b).
Appendix F Constraint Satisfaction Problem and Soft Constraint Satisfaction Problems